Hi everyone, does any one would be able to help me with this question pease?

Find the V (value) of a MNC based on the following information.

t=1 t=2

Exp. CF from local market 200 200

Exp. CF from a foreign market 300 . 500

Exp. Exchange rate (S) 0.6 . 0.7

Required rate of return (k) . 0.2   0.2
Answer #1

Value at Year 1 =

Cash Flow from Local Market = 200

Cash Flow from Foreign Market = 300

Equivalent Cash Flow from Foreign Market in Local Market = 300 * 0.6 = 180

Total Cash Flow = 200+180 = 380

Value of MNC = Cash Flow / Required Rate of return

= 380/0.2

= 1900

Value at Year 2 =

Cash Flow from Local Market = 200

Cash Flow from Foreign Market = 500

Equivalent Cash Flow from Foreign Market in Local Market = 500 * 0.7 = 350

Total Cash Flow = 200+350 = 550

Value of MNC = Cash Flow / Required Rate of return

= 550/0.2

= 2750